OpenStudy (anonymous):

Find the 99th derivative

OpenStudy (anonymous):

if you have a function \[f(x) = \frac{ \cos^{2}x- \sin^{2}x }{ secxcscx }\]

OpenStudy (anonymous):

evaluate it at pi/12

OpenStudy (primeralph):